# Robust Design and Evaluation of

### v5

@@ C.4.3 Bounds on overall performance
Like in our analysis of nonparametric outcome bounds in Section 4.2, the bounds on overall performance are again the expectation of some known function of the data and nuisance parameters $\eta=(\lambda_{z}(\cdot),\kappa_{z}(\cdot),\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot))$. We therefore again randomly split the data into $K$ disjoint folds. For each fold $k$, we estimate the nuisance functions $\hat{\eta}_{-k}$ using all observations not in the $k$-th fold and construct
@@
Proposition C.6.
Define $R_{2,n}^{k}=\|\hat{\lambda}_{z,-k}(\cdot)-\lambda_{z}(\cdot)\|\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot)-\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot)\|$ and $R_{3,n}^{k}=\|\hat{\kappa}_{z,-k}(\cdot)-\kappa_{z}(\cdot)\|\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot)-\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot)\|$. Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}\{\mathbb{P}(Z_{i}=z\mid X_{i})\geq\delta\}=1$; (iii) there exists $\epsilon>0$ such that {^$\mathbb{P}\{\hat{\mathbb{P}}_{-k}(Z_{i}=z\mid X_{i})\geq\epsilon\}=1$ for all folds $k$; and (iv) $\|\hat{\lambda}_{z}(\cdot)-\lambda_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, $\|\hat{\kappa}_{z}(\cdot)-\kappa_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, and ^$\|\hat{\mathbb{P}}_{-k}(Z_{i}=z\mid X_{i})-\mathbb{P}(Z_{i}=z\mid X_{i})\|=o_{\mathbb{P}}(1)$ for all folds $k$. Then,
@@
C.4.4 Smooth approximations to intersection bounds
@@
$\underline{\delta}(x;\eta)=g_{\alpha}\left(\lambda_{z_{1}}(x)-\mu_{1}(x),\ldots,\lambda_{z_{p}}(x)-\mu_{1}(x)\right)\mbox{ and }$
@@
are the uncentered efficient influence functions for $\mathbb{E}[\lambda_{z}(X_{i})]$ and $\mathbb{E}[\kappa_{z}(X_{i})]$ respectively. Furthermore,
@@
are the uncentered efficient influence functions for $\mathbb{E}[\underline{\delta}(X_{i};\eta)]$, $\mathbb{E}[\overline{\delta}(X_{i};\eta)]$ respectively, by Theorem 3 in LevisEtAl(23)-CovariateAssistedIV and standard calculations involving influence functions. With this in hand, we then can directly extend our estimators for the bounds on overall performance, positive class performance, and the conditional probability by substituting in the appropriate uncentered efficient influence functions for the upper bounding and lower bounding functions.
C.5 Identification and estimation of predictive disparities
Consider a binary sensitive attribute $G_{i}\in\{0,1\}$ (e.g., ethnicity, gender, race, etc.), define the overall disparity of $s(\cdot)$ as $\mbox{disp}(s;\beta):=\mbox{perf}_{1}(s;\beta)-\mbox{perf}_{0}(s;\beta)$ for $\mbox{perf}_{g}(s;\beta):=\mathbb{E}[\beta_{0}(X_{i},s)+\beta_{1}(X_{i};s)Y_{i}^{*}\mid G_{i}=g]$ the overall performance on group $G_{i}=g$. The positive class and negative class disparities $\mbox{disp}_{+}(s;\beta)$ and $\mbox{disp}_{-}(s;\beta)$ are defined analogously.
In this section, we show that predictive disparities as defined in Section 2.1 are partially identified under Equation (6). We then propose estimators for the bounds on predictive disparities, extending our analysis of overall and positive class performance.
C.5.1 Positive class predictive disparities
We provide non-sharp bounds for the positive class disparity since the positive class disparity is the difference of two linear-fractional functions. Define $\mathcal{H}(\mbox{disp}_{+}(s;\beta))$ to be the set of all positive class disparities that are consistent with Equation (6).
Lemma C.5.
Proof.
Observe that, for $g\in\{0,1\}$ and $\delta(\cdot)\in\Delta$,
and
We can therefore construct estimators for the non-sharp bounds on positive class disparities, extending our estimators from the main text. To do so, we estimate the bounds on group-specific positive class performance $\overline{\mbox{perf}}_{+,g}(s;\beta)$ and $\underline{\mbox{perf}}_{+,g}(s;\beta)$. For simplicity, we again develop the estimators assuming $\mathbb{P}(G_{i}=1)$ is known and for nonparametric outcome bounds.
We once again split the data into three folds. We construct nuisance function estimate $\hat{\eta}$ using only the first fold. On the second fold, we construct the pseudo-outcomes $\phi_{\mu,i}(\hat{\eta})$ and regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified nonparametric outcome procedure, yielding $\widehat{\delta}(X_{i})$. On the third fold, we construct an estimate of the upper bound for group $g$ by solving
$\widehat{\overline{\mbox{perf}}}_{+,g}(s;\beta):=\max_{\tilde{\delta}\in\widehat{\Delta}_{n,k}}\frac{\mathbb{E}_{n}[1\{G_{i}=g\}\left(\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}\right)]}{\mathbb{E}_{n}[1\{G_{i}=g\}\left(\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}\right)]}$ (15)
for $\widehat{\Delta}_{n,k}:=\left\{(\underline{\Gamma}-1)(1-D_{i})\phi_{\mu,i}(\hat{\eta})\leq\widetilde{\delta}_{i}\leq(\overline{\Gamma}-1)(1-D_{i})\phi_{\mu,i}(\hat{\eta})\right\}$. The estimator $\widehat{\underline{\mbox{perf}}}_{+,g}(s;\beta)$ is defined analogously. We then estimate the bounds on positive class disparities by $\widehat{\overline{\mbox{disp}}}_{+}(s;\beta)=\widehat{\overline{\mbox{perf}}}_{+,1}(s;\beta)-\widehat{\underline{\mbox{perf}}}_{+,1}(s;\beta)$ and $\widehat{\underline{\mbox{disp}}}_{+}(s;\beta)=\widehat{\underline{\mbox{perf}}}_{+,1}(s;\beta)-\widehat{\overline{\mbox{perf}}}_{+,1}(s;\beta)$. By the same arguments as the proof of Proposition 4.1 and Corollary 4.1, we can derive the rate of convergence of our estimators for positive class disparities.
C.5.2 Overall predictive disparities
For $g\in\{0,1\}$, observe that
$\mbox{perf}_{g}(s;\beta)=P(G_{i}=g)^{-1}\mathbb{E}[\beta_{0,i}1\{G_{i}=g\}+\beta_{1,i}1\{G_{i}=g\}\mu_{1}(X_{i})+\beta_{1,i}1\{G_{i}=g\}\pi_{0}(X_{i})\delta(X_{i})].$
where $\alpha_{g}=P(G_{i}=g)$. Therefore the overall disparity $\mbox{disp}(s;\beta)$ can be equivalently written as
$\alpha_{1}^{-1}\mathbb{E}[\beta_{0,i}G_{i}+\beta_{1,i}G_{i}\mu_{1}(X_{i})+\beta_{1,i}G_{i}\pi_{0}(X_{i})\delta(X_{i})]-\alpha_{0}^{-1}\mathbb{E}[\beta_{0,i}(1-G_{i})+\beta_{1,i}(1-G_{i})\mu_{1}(X_{i})+\beta_{1,i}(1-G_{i})\pi_{0}(X_{i})\delta(X_{i})].$
Define $\mathcal{H}(\mbox{disp}(s;\beta))$ to be the set of all overall predictive disparities that are consistent with Equation (6), and further let $\beta_{0,i}^{g}:=\beta_{0,i}/P(G_{i}=g)$, $\beta_{1,i}^{g}:=\beta_{1,i}/P(G_{i}=g)$ for $g\in\{0,1\}$, and $\tilde{\beta}_{0,i}:=G_{i}\beta_{0,i}^{1}-(1-G_{i})\beta_{0,i}^{0}$, $\tilde{\beta}_{1,i}:=G_{i}\beta_{1,i}^{1}-(1-G_{i})\beta_{1,i}^{0}$.
Lemma C.6.
where
Proof.
To prove the result, we rewrite $\mbox{disp}(s;\beta)$ as
using the definitions of $\tilde{\beta}_{0,i},\tilde{\beta}_{1,i}$. As in the proof of Lemma 2.1, it then follows immediately that $\mathcal{H}(\mbox{disp}(s;\beta))$ equals the closed interval
$[\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})\left(1\{\tilde{\beta}_{1,i}\geq 0\}\underline{\delta}_{i}+1\{\tilde{\beta}_{1,i}<0\}\overline{\delta}_{i}\right),$
The result then follows by noticing that
Since the bounds on overall disparities are linear functionals of known functions of the data and identified nuisance parameters, we construct debiased estimators for the bounds on overall disparities as in the main text. For simplicity, we develop the estimators assuming $\mathbb{P}(G_{i}=1)$ is known and for nonparametric outcome bounds, but they can be easily extended to the case where this is estimated and for alternative choices of bounding functions.
We make use of $K$-fold cross-fitting. For each fold $k$, we construct estimators of the nuisance functions $\hat{\eta}_{-k}=\left(\hat{\pi}_{1,-k},\hat{\mu}_{1,-k}\right)$ using only the sample of observations not in the $k$-th fold. For each observation in the $k$-th fold, we construct
for $\hat{\sigma}_{i,11}=(\overline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\overline{\mbox{disp}}}(s;\beta))^{2}$, $\hat{\sigma}_{i,12}=(\overline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\overline{\mbox{disp}}}(s;\beta))(\underline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\underline{\mbox{disp}}}(s;\beta))$, and $\hat{\sigma}_{i,22}=(\underline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\underline{\mbox{disp}}}(s;\beta))^{2}$. By the same arguments as the proof of Proposition 4.2, we can derive the rate of convergence of our proposed estimators and provide conditions under which they are jointly asymptotically normal.
Appendix D Connections to alternative sensitivity analysis models
As mentioned in Section 2.3.1 of the main text, nonparametric outcome bounds in our setting with a binary outcome are related to popular sensitivity analysis models for unobserved confounding in causal inference.
Marginal sensitivity model:
A recently popular sensitivity analysis model is the marginal sensitivity model (MSM). The MSM specifies that $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies, for some $\Lambda\geq 1$,
$\Lambda^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid X_{i},Y_{i}^{*})}{\mathbb{P}(D_{i}=0\mid X_{i},Y_{i}^{*})}\frac{\mathbb{P}(D_{i}=0\mid X_{i})}{\mathbb{P}(D_{i}=1\mid X_{i})}\leq\Lambda$ (18)
with probability one. The MSM nests unconfoundedness by setting $\Lambda=1$. The MSM was originally proposed by Tan(06) and has since received substantial attention among researchers (e.g., see ZhaoSmallBhattacharya(19); KallusMaoZhou(18)-IntervalEstimation; DornGuo(21); DornGuoKallus(21); KallusZhou(21); JinRenCandes(21)).
Proposition D.1.
@@ Appendix D Connections to alternative sensitivity analysis models
    Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies the MSM (18) for some $\Lambda\geq 1$. Then, $\mathbb{P}(\cdot)$ satisfies Equation (6) with $\underline{\delta}(x;\eta)=(\Lambda^{-1}-1)\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=(\Lambda-1)\mu_{1}(x)$.
@@
    Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies Equation (6) with nonparametric outcome bounds for some $\underline{\Gamma},\overline{\Gamma}>0$. Then, $\mathbb{P}(\cdot)$ satisfies
@@
Consider the second claim. Nonparametric outcome bounds implies
$\overline{\Gamma}^{-1}\leq\frac{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)}\leq\underline{\Gamma}^{-1}.$
But, by Bayes’ rule, $\frac{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)}=\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=1)}$, and so
$\overline{\Gamma}^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=1)}\leq\underline{\Gamma}^{-1}$
holds. ∎
This first result relates to Proposition 3 in DornGuoKallus(21), which establishes that the MSM implies a bound on $\mathbb{E}[Y_{i}^{*}\mid D_{i}=0,X_{i}]$ via the solution to a conditional value-at-risk problem. For the binary outcome setting, we show the MSM directly implies a bound on $\mathbb{E}[Y_{i}^{*}\mid D_{i}=0,X_{i}]$ via Bayes’ rule. Our second result establishes a partial converse, showing that Equation (6) implies an MSM-like bound for sample-selection models (e.g., our running credit lending and pretrial release examples).
Rosenbaum’s sensitivity model:
Another popular sensitivity analysis model is Rosenbaum’s $\Gamma$-sensitivity analysis model (e.g., Rosenbaum(87); Rosenbaum(02)). The $\Gamma$-sensitivity analysis model specifies that for some $\Gamma\geq 1$, $(X_{i},D_{i},Y_{i}(0),Y_{i}(1))\sim\mathbb{P}(\cdot)$ satisfies

### v6

@@ D.2.2 Bounds on Overall Performance
Like in Section 4.1, the bounds are again linear functionals of known functions of the data and identified nuisance parameters. We construct debiased estimators using standard arguments.
Estimation Procedure for Single-Instrument Bounds:
We randomly split the data into $K$ disjoint folds. For each fold $k$, we estimate the nuisance functions $\hat{\eta}_{-k}$ using all observations not in the $k$-th fold and construct
@@
Proposition D.4.
Define $R_{2,n}^{k}=\|\hat{\lambda}_{z,-k}(\cdot)-\lambda_{z}(\cdot)\|\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot)-\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot)\|$ and $R_{3,n}^{k}=\|\hat{\kappa}_{z,-k}(\cdot)-\kappa_{z}(\cdot)\|\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot)-\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot)\|$. Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}\{\mathbb{P}(Z_{i}=z\mid X_{i})\geq\delta\}=1$; (iii) there exists $\epsilon>0$ such that $\mathbb{P}\{\hat{\mathbb{P}}_{-k}(Z_{i}=z\mid X_{i})\geq\epsilon\}=1$ for all folds $k$; and (iv) $\|\hat{\lambda}_{z}(\cdot)-\lambda_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, $\|\hat{\kappa}_{z}(\cdot)-\kappa_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, and $\|\hat{\mathbb{P}}_{-k}(Z_{i}=z\mid X_{i})-\mathbb{P}(Z_{i}=z\mid X_{i})\|=o_{\mathbb{P}}(1)$ for all folds $k$. Then,
@@
$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)\xrightarrow{d}N\left(0,\Sigma\right)$
D.2.3 Bounds on Positive Class Performance
We analogously estimate the instrumental variable bounds on positive class performance by solving sample linear fractional programs as in Section 4.2 of the main text. All that needs to be modified is the choice estimator for the bounding functions. As an illustration, we sketch out how the sample linear fractional program can be combined with our pseudo-regression procedure discussed earlier in Appendix D.2.1.
Estimation Procedure under Single-Instrument Bounds:
@@ D.2.3 Bounds on Positive Class Performance
where now $\widehat{\Delta}_{n}=\{(1-D_{i})\widehat{\underline{\delta}}(X_{i})\leq\tilde{\delta}_{i}\leq(1-D_{i})\widehat{\overline{\delta}}(X_{i})\}$. Proposition 4.2 can be applied, and this estimator has the same partial double robustness property. The error of this estimator again depends on the root mean square error of the estimated bounding functions. If the nonparametric regression procedure further satisfies the $L_{2}(\mathbb{P})$-stability condition, then we can further control these errors using the results in Appendix D.2.1.
D.2.4 Smooth Approximations to Intersection Bounds
@@
[…]
@@
are the influence functions for $\mathbb{E}[\lambda_{z}(X_{i})]$ and $\mathbb{E}[\kappa_{z}(X_{i})]$ respectively. Furthermore,
@@
are the influence functions for $\mathbb{E}[\underline{\delta}(X_{i};\eta)]$, $\mathbb{E}[\overline{\delta}(X_{i};\eta)]$ respectively, by Theorem 3 in LevisEtAl(23)-CovariateAssistedIV and standard calculations involving influence functions. With this in hand, we then extend our estimators for the bounds on overall performance, positive class performance, and the conditional likelihood by substituting in the appropriate influence functions for the upper bounding and lower bounding functions.
Alternative Approach under Margin Conditions.
An alternative is to invoke a margin condition as in LevisEtAl(23)-CovariateAssistedIV and semenova2023adaptive. The margin condition assumes that, for each $x\in\mathcal{X}$, the maximal and minimal values in the intersection bounds are sufficiently well-separated from the remaining instrument values. Under such margin conditions, LevisEtAl(23)-CovariateAssistedIV and semenova2023adaptive show that one can construct estimators directly targeting the averages of intersection bounds. For our overall performance bounds, these results apply directly. Extending this approach to estimate the conditional likelihood bounds and the positive class performance bounds is non-trivial. Developing these extensions is an interesting direction for future work.
Appendix E Connections to Sensitivity Analysis Models
A substantial literature in causal inference develops sensitivity analysis frameworks for assessing robustness to unobserved confounding. While these frameworks are often presented in treatment effect settings, they can be translated to our selective labels context. This appendix establishes formal connections between our framework—specifically Assumption 2.1 with observed outcome bounds discussed in Section 2.3.1 of the main text—and several influential sensitivity analysis models.
This is valuable for three reasons. First, pedagogically, researchers familiar with sensitivity analysis in causal inference may find it easier to interpret observed outcome bounds by relating them to established frameworks. Second, these connections show that our estimation results extend to settings where researchers invoke these alternative sensitivity models. Rather than developing separate estimation procedures for each model, researchers could apply our results by translating their sensitivity assumptions into the corresponding values of $\underline{\Gamma},\overline{\Gamma}$ (although the resulting bounds may not be sharp). Finally, most sensitivity analysis models in causal inference place restrictions on selection propensities—how unobservables affect the probability of treatment/selection. In contrast, Assumption 2.1 with observed outcome bounds restricts outcome differences—how outcomes differ between selected and unselected units conditional on observables. While conceptually distinct, Bayes’ rule establishes precise mathematical connections between these two approaches.
E.1 Marginal Sensitivity Model
The marginal sensitivity model (MSM) has received substantial recent attention in the causal inference literature (Tan(06); ZhaoSmallBhattacharya(19); KallusMaoZhou(18)-IntervalEstimation; DornGuo(21); DornGuoKallus(21); KallusZhou(21); JinRenCandes(21)). The MSM specifies that the selection mechanism satisfies a multiplicative bound on odds ratios: that is, for some $\Lambda\geq 1$,
$\Lambda^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid X_{i},Y_{i}^{*})}{\mathbb{P}(D_{i}=0\mid X_{i},Y_{i}^{*})}\frac{\mathbb{P}(D_{i}=0\mid X_{i})}{\mathbb{P}(D_{i}=1\mid X_{i})}\leq\Lambda\mbox{ holds with probability }1.$ (15)
The MSM bounds the ratio of conditional-to-marginal odds of selection. When $\Lambda=1$, this reduces to $\mathbb{P}(D_{i}=1\mid X_{i},Y^{*}_{i})=\mathbb{P}(D_{i}=1\mid X_{i})$ i.e., unconfoundedness. Larger values of $\Lambda$ allow greater departure from unconfoundedness. In other words, the parameter $\Lambda$ directly quantifies the strength of confounding on the odds ratio scale.
We can relate the MSM to Assumption 2.1 under observed outcome bounds via Bayes’ rule.
Proposition E.1.
@@ E.1 Marginal Sensitivity Model
    Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies the MSM (15) for some $\Lambda\geq 1$. Then, $\mathbb{P}(\cdot)$ satisfies Assumption 2.1 with $\underline{\delta}(x;\eta)=(\Lambda^{-1}-1)\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=(\Lambda-1)\mu_{1}(x)$.
@@
    Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies Assumption 2.1 with observed outcome bounds for some $\underline{\Gamma},\overline{\Gamma}>0$. Then, $\mathbb{P}(\cdot)$ satisfies
@@
Consider the second claim. Observed outcome bounds implies $\overline{\Gamma}^{-1}\leq\frac{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)}\leq\underline{\Gamma}^{-1}$. But, by Bayes’ rule, $\frac{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)}=\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=1)}$, and so $\overline{\Gamma}^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=1)}\leq\underline{\Gamma}^{-1}$ holds. ∎
Proposition E.1 shows that any analysis conducted under the MSM with parameter $\Lambda$ can be conducted using observed outcome bounds with $\underline{\Gamma}=\Lambda^{-1}$ and $\overline{\Gamma}=\Lambda$. In the opposite direction, observed outcome bounds imply a specific MSM-style restriction on selection propensities, but only for the positive class $Y_{i}^{*}=1$.
E.2 Rosenbaum’s $\Gamma$-Sensitivity Model
Rosenbaum’s $\Gamma$-sensitivity model (e.g., Rosenbaum(87); Rosenbaum(02)) is perhaps the most widely used framework for sensitivity analysis in observational studies. It bounds how much unobservables can differentially affect selection propensities across different outcome values: specifically, for some $\Gamma\geq 1$, $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies